Grade 11Mechanics

a bomb explodies into two fragments of masses 2kg and 1kg. The total KE of the frangments is 3×10to yhe power 4 joules.Find the KE of the bigger frangment And momentum of the smaller fragment?

sachin sharma
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
From law of conservation of momentum
 
m1 * v1 = m2 * v2
v1/v2 = ½
Total K.E. = 3 * 104
½* m1 * (v1)² + ½ * m2 *(v2)² = 3 *104
3 (v1)² = 3 *104
(v1) = 10² = 100 m/sec
Now
K.E. of bigger fragment = 1/2 * 2 * 100² = 10000 J
Momentum of smaller fragment = m2 * v2
 = 1 * 2 * 100 = 200 kg - m/sec
 
Regards
Arun (askIITians forum expert)